Two point charges, +3.72 µC and -5.75 µC, are separated by 1.25 m. What is the electric potential midway between them?

I am doing U(total)=U1+U2 and am getting the wrong answer. I am not sure where I am going wrong.

U1=2kq1/d

U2=2kq2/d
Answer #1

U1 = kq1 / (d/2) = 2k q1 / d

U1 = 2*9*10^9*3.72*10^-6/1.25

U1 = 53568


U2 = kq2 / (d/2) = 2k q2 / d

U2 =  2*9*10^9*-5.75*10^-6/1.25

U2 = 82800


so U total = U1 + U2 = (2k/d) (q1 + q2) , d = 1.3 m


U total =   2*9*10^9*/1.25(3.72*10^-6+-5.75*10^-6)


U total = -0.308016 V